Bellingham Company produces a product that requires 6 standard hours per unit at a standard hourly rate of $17.00 per hour. If 4,900 units required 28,200 hours at an hourly rate of $17.68 per hour, what is the direct labor (a) rate variance, (b) time variance, and (c) total direct labor cost variance? Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

a. Direct labor rate variance $
b. Direct labor time variance $
c. Total direct labor cost variance $
Answer #1

Labor rate variance = (SR-AR) *AH

= (17-17.68)*28200

= 19176 Unfavorable

Labor time variance = (SH-AH) *SR

= (4900*6 - 28200)*17

= - 20,400 Favourable

Labor cost variance = 19176U + 20400F

= - 1224 Favourable